For a 20 inch square base plate of A36 steel supporting a W12 x 50 column,determine the axial compression force from the column Pu.The base plate is on the full area(A1=A2)of a concrete pedestal (fc=5ksi).The baseplate thickness t= 15/8 inch.
1
Expert's answer
2020-06-11T15:11:50-0400

Ψ

\delta\Psi

δΨ

 Since:

φ_"c" p_p=0.85φ_"c" F_c^,A_1 √(A_2⁄A_1 )≤φ_"c" 1.7F_c^,A_1

0.85P_u=0.85×0.85×(12×50)×400√(400⁄400)≤0.85×1.7×(12×50)400

0.85P_u=173400≤346800

P_u=173400/0.85≤346800

P_u=204000≤346800

P_u=204000/346800

P_u=0.5882N/〖In〗^2